OpenStudy (anonymous):

let f(x)=4+5x and g(x)=2x-1 what is g(f(x)) ?

OpenStudy (anonymous):

for this you set it up like g(f(x)) = 2(4+5x)-1 = 8+10x-1 = 10x-7

OpenStudy (anonymous):

thank you!!!!!!!